OpenStudy (anonymous):

(-16b^3) (1/20b^3) (-25b^2) simplify

OpenStudy (king):

20b^2

OpenStudy (king):

see in the 1st 2 expression b^3 get cancelled and the u get16*25/20=20 so answer is 20b^2

OpenStudy (diyadiya):

\[(-16{\cancel b^3}) (\frac{1}{20{\cancel b^3}}) (-25b^2) \] \[\frac{-16}{20} \times -25b^2= \frac{-16}{4} \times -5 b^2= -4 \times -5 b^2 =20b^2\]

OpenStudy (anonymous):

its not a 20b^2

OpenStudy (diyadiya):

it is 20b^2 see the solution :)